OpenStudy (anonymous):

A sandbag was thrown downward from a building. The function f(t) = -16t2 - 32t + 384 shows the height f(t), in feet, of the sandbag after t seconds: Part A: Factor the function f(t) and use the factors to interpret the meaning of the x-intercept of the function. Part B: Complete the square of the expression for f(x) to determine the vertex of the graph of f(x). Would this be a maximum or minimum on the graph? Part C: Use your answer in part B to determine the axis of symmetry for f(x)? @welshfella

OpenStudy (welshfella):

first simplify the function by taking out -16 so y = -16(t^2 + 2t - 24 )

OpenStudy (welshfella):

can you factor the quadratic in the parentheses?

OpenStudy (anonymous):

yes hang on one sec please

OpenStudy (anonymous):

f(t)=−16(t+6)(t−4)

OpenStudy (anonymous):

So thats it for part A?

OpenStudy (anonymous):

wait what would be the meaning???

OpenStudy (welshfella):

right now f(t) = the height of the sandbag at t seconds so it you put f(t) = 0 what do the values of t signify?

OpenStudy (anonymous):

t is seconds?

OpenStudy (welshfella):

- there will be 2 x-intercepts

OpenStudy (welshfella):

t is seconds yes

OpenStudy (anonymous):

wait im lost sorry

OpenStudy (anonymous):

ok so Part A would be f(t)=−16(t+6)(t−4) where t is equaled to seconds?

OpenStudy (welshfella):

f(t) = -16(t + 6)( t - 4) = 0 find values of t

OpenStudy (welshfella):

- the x intercepts are values of t when f(t) the height = 0.

OpenStudy (anonymous):

So the x-intercepts or t are equaled to -6 and 4?

OpenStudy (welshfella):

right - now you neglect -6 The value t = 4 means the time when height of the sandbag from the ground = 0. That is when it lands.

OpenStudy (anonymous):

So Part A is the factored function: f(t)=−16(t+6)(t−4) where the x-intercepts or t represents the time in seconds that the sandbag was thrown?

OpenStudy (anonymous):

and the x-intercept is (4,0) which represents the time in seconds that the sandbag was thrown?

OpenStudy (anonymous):

to where it lands

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

for Part b the completing of the square is: (t+1)^2=25 right

OpenStudy (welshfella):

yes t = 0 is the time when in is thrown and t = 4 = when it lands

OpenStudy (welshfella):

don't forget the -16

OpenStudy (anonymous):

ok so part a is complete!

OpenStudy (anonymous):

Wait the -16 for what??

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

part a is complete yes the function is -16(t^2 + 2t - 24)

OpenStudy (anonymous):

the function is f(t)=−16(t+6)(t−4 For Part B: complete the square: (t+1)^2=25

OpenStudy (welshfella):

-16(t^2 + 2t - 24) = -16 [ (t + 1)^2 - 15]

OpenStudy (anonymous):

now I have to find the vertex and decide whether it is a maximum or minimum

OpenStudy (welshfella):

= -16 [ (t + 1)^2 - 25]

OpenStudy (welshfella):

yes it will be a maximum because the coefficient of t^2 is negative

OpenStudy (anonymous):

ok so (t+1)^2=25 is a maximum with the vertex or 400??

OpenStudy (anonymous):

of^^^

OpenStudy (welshfella):

excuse me i 'll be back in a minute or two...

OpenStudy (anonymous):

ok

OpenStudy (welshfella):

right the maximum point on the graph is when t+1 = 0 that s is t = -1 and f(t0 will equal 400 (-1,400)

OpenStudy (welshfella):

the axis of symmetry is the vertical line passing through the point (-1,400) which is x = -1

OpenStudy (welshfella):

* t = -1

OpenStudy (anonymous):

ok I understand thanks so much! Is that everything??

OpenStudy (welshfella):

oh they talk about f(x) in parts b and c which makes more sense than t as you cant have negative times! so - change all those references to t to x.

OpenStudy (anonymous):

ok thank you! @welshfella I may need more help on other questions Ill tag you!!!!!!!!

OpenStudy (welshfella):

ok yw